A 32 g bullet is fired horizontally from a rifle a
Answer #1

a )

F is the average force excerted by water on bullet

F X d X cos180 = 0.5 X m ( V22 - V12 )

F X 0.12 X ( -1 ) = 0.5 X 0.032 (02 - 3802 )

F X ( -0.12 ) = -2310.4

F = 19253.33 N

b )

the acceleration of the bullet a = - F / m

= - 19253.33 / 0.032

a = - 601666.66 m/s2

the time is = ( V2 - V1 ) / a

= ( 0 - 380 ) / ( - 601666.66 )

= 0.631 ms

and t = 2 h / g

t = ( 2 X 2 / 9. 8

t = 0.6388 sec

the total time is t = 0.631 X 10-3 + 0.6388

T = 0.6394 sec